# Scalable Causal Discovery from Recursive Nonlinear Data via Truncated Basis Function Scores and Tests

### v1

@@ 1 Introduction
We propose two basis–expansion tools for recursive nonlinear Structural Equation Models (rSEMs): the Basis Function BIC (BF-BIC) and the Basis Function Likelihood Ratio Test (BF-LRT). Both use truncated orthogonal polynomial expansions of continuous variables to represent additive nonlinear dependencies in finite-dimensional linear spaces. BF-BIC is a scoring criterion for score-based search, while BF-LRT serves as an approximate conditional independence test for constraint-based search.
@@
Although the working assumption is additive nonlinearity (each variable modeled as a nonlinear sum of functions of its parents), the methods remain robust under moderate interactions and generalize formally to the post–nonlinear (PNL) setting via an invertible reparameterization (Section B.6 of the appendix). Thus, the additive formulation is not restrictive in practice.
As a real-data demonstration, we apply BF-BIC and BF-LRT within the FCIT algorithm to the Algerian Forest Fire dataset. This hybrid search recovers plausible nonlinear causal structure, including known relationships such as $\text{FWI}\rightarrow\text{Fire}$, while supporting latent-variable discovery via Partial Ancestral Graphs (PAGs). This illustrates the compatibility of our methods with latent-variable causal discovery and their utility for scientific modeling in domains with both deterministic indices and hidden confounders.
@@ 4 Basis Function BIC and Basis Function LRT
In the general case—when the structural function contains interactions or non-additive components—BF-BIC still provides a flexible and scalable approximation. Although the residuals may no longer be exponential-family distributed, empirical results indicate that BF-BIC performs well in practice, offering a heuristic but effective model selection criterion.
@@
The theoretical justification for BF-LRT is provided in Appendix B. As with BF-BIC, theoretical guarantees apply in the additive case under exponential-family assumptions. In more general nonlinear settings, BF-LRT remains a scalable and effective heuristic for independence testing.
@@
Other approaches to nonlinear causal discovery include the use of post-nonlinear models (Zhang and Hyvärinen, 2009a, ), general non-Gaussian SEMs (Zhang and Hyvärinen,, 2010), and deep generative models for causal discovery (Zhang et al.,, 2016). These models typically emphasize functional identifiability under minimal assumptions, often at the cost of scalability. In contrast, our approach focuses on scalability through a truncated basis representation, offering a practical tradeoff between flexibility and interpretability.
@@ 5 Using BF-BIC and BF-LRT for Inferring CPDAG Models
BF-BIC and BF-LRT were developed to support causal search algorithms for estimating nonlinear CPDAG models—that is, Markov equivalence classes of DAGs without latent confounders. While both methods are theoretically consistent under additive structural functions, they may also be used heuristically in broader nonlinear settings. In practice, they can serve as plug-in tools wherever a nonlinear score or conditional independence test is required.
@@ 6 Simulation Comparisons
We compare BOSS/BF-BIC, PC-Max/BF-LRT, and PC-Max/KCI (Zhang et al.,, 2012) and PC-Max/RCIT (Strobl et al.,, 2019) in smaller-scale simulations. We use KCI and RCIT from the causal-learn package (Zheng et al.,, 2024). KCI uses kernel reasoning to ensure that $cov(f(X),g(Y))=0$ for all $f$ and $g$ for $X{\;\perp\!\!\!\perp\;}Y$, for the unconditional case, and uses similar reasoning for the conditional case. Note that KCI in causal-learn is fully parallelized, with all matrix calculations optimized for multi-core execution. RCIT chooses random Fourier features to model non-linear relationships. Like BF-BIC, these methods consider functions of $X$ and $Y$, though in the context of the theorem that $cov(f(X),g(Y))$ being zero for all $f$ and $g$ if and only if $X{\;\perp\!\!\!\perp\;}Y$. Fourier features are random sin or cosine functions of $X$ or $Y$. Because these implementations are slower than the BF methods for large $N$ and $p$, we include them only for the continuous 10 node average degree 2 cases. [^3] All simulations are performed on a MacBook Pro M2 with 64 GB of RAM and 10 cores. RCIT scales fairly well though is dominated by the basis function methods. This may be partially due to its being implemented in Python, which is a slower language than Java. For speed, though, an implementation in C would be advantageous; we may pursue this in future work.
The CAM method proposed by Bühlmann et al., (2014) is directly relevant to our setting, as it assumes a sparse, additive structural equation model with Gaussian noise. [^4] For a feature comparison of CAM, additive noise models, and our BF-BIC and BF-LRT methods, please see Table 1 in the Appendix. Although we had initially planned to include CAM in our benchmarking suite, we encountered technical challenges: the official R package for CAM has been removed from CRAN and is no longer actively maintained, and the provided implementation could not be easily integrated into our Python-based pipeline via rpy2. While it is possible that more effort on our part could have resolved this, we were ultimately unable to incorporate CAM into our large-scale simulation framework in a fully automated and reproducible manner.
To provide a meaningful point of comparison, however, we replicated the experimental setup from Figure 3 of the CAM paper, generating random DAGs with 10 nodes, average in-degree 2, and 1000 samples, using additive nonlinear functions as in Peters et al., (2014) with Gaussian noise. On this benchmark, BOSS with BF-BIC achieved an average SHD of 5.40 (±3.85), adjacency precision 1.00 (±0.00), and adjacency recall 0.90 (±0.06) over 10 runs. These results indicate strong performance in the same setting used to validate CAM. Our script for this comparison has been included on our GitHub site.
@@ 9 Conclusion
In our real-data analysis of the Algerian Forest Fire dataset, we applied the FCIT algorithm—an accurate hybrid of test and score supporting latent-variable search—and recovered plausible causal structures in the form of Partial Ancestral Graphs (PAGs). In the full model, which included deterministic Canadian Fire System indices, FCIT correctly identified the expected edge $\text{FWI}\rightarrow\text{Fire}$ along with much of the computational structure underlying fire risk indices. The measured-only model also recovered interpretable predictors such as RH and Temperature, reinforcing the reliability of our approach. These results suggest that combining FCIT with BF-based scores and tests yields informative models even in the presence of nonlinearities and latent confounding.
@@
Thanks to Peter Spirtes and Clark Glymour for helpful comments on earlier drafts.
BA was supported by the US National Institutes of Health under the Comorbidity: Substance Use Disorders and Other Psychiatric Conditions Training Program T32DA037183. JR was supported by the US Department of Defense under Contract Number FA8702-15-D-0002 with Carnegie Mellon University for the operation of the Software Engineering Institute. The content of this paper is solely the responsibility of the authors and does not necessarily represent the official views of these funding agencies. The authors have no conflict of interest to report.

### v2

@@ 1 Introduction
We propose two basis–expansion tools for recursive nonlinear Structural Equation Models (rSEMs): the Basis Function BIC (BF-BIC) and the Basis Function Likelihood Ratio Test (BF-LRT). Both use truncated orthogonal polynomial expansions of continuous variables to represent additive nonlinear dependencies in finite-dimensional linear spaces. BF-BIC is a scoring criterion for score-based search, while BF-LRT serves as an approximate conditional independence test for constraint-based search. For methods such as FCIT (Ramsey and Andrews,, 2025) that use both a score and a test, the two methods may be used in tandem.
@@
Although the working assumption is additive nonlinearity (each variable modeled as a sum of nonlinear functions of its parents), the methods generalize formally to the post–nonlinear (PNL) setting via an invertible reparameterization (Section B.6 of the appendix). Thus, the additive formulation is not restrictive in practice.
As a real-data demonstration, we apply BF-BIC and BF-LRT within the FCIT algorithm to the Algerian Forest Fire dataset. This hybrid search recovers plausible nonlinear causal structure, including known relationships such as $\text{FWI}\rightarrow\text{Fire}$, while supporting latent-variable discovery via Partial Ancestral Graphs (PAGs). We also note a recent astrophysics note (Desmond and Ramsey,, 2025) using FCIT (Ramsey and Andrews,, 2025) with both BF-BIC and BF-LRT to model galaxy data. These applications illustrate the compatibility of our methods with latent-variable causal discovery and their utility for scientific modeling in domains with both deterministic indices and hidden confounders.
@@ 4 Basis Function BIC and Basis Function LRT
In the general case—when the structural functions include interactions or other non-additive components—BF-BIC remains a flexible and scalable approximation. Although the residuals may then deviate from exponential-family form, the framework can be extended to the post-nonlinear setting by applying an inverse mapping to the additive nonlinear component of each structural equation.
@@
The theoretical justification for BF-LRT is provided in Appendix B. As with BF-BIC, the formal guarantees hold in the additive case under exponential-family assumptions, and extend to post-nonlinear models formed by applying invertible transformations to the additive components.
@@
Other approaches to nonlinear causal discovery include general non-Gaussian SEMs (Zhang and Hyvärinen,, 2010) and deep generative models for causal discovery (Zhang et al.,, 2016). These models typically emphasize functional identifiability under minimal assumptions, often at the cost of scalability. In contrast, our approach focuses on scalability through a truncated basis representation, offering a practical tradeoff between flexibility and interpretability.
@@ 5 Using BF-BIC and BF-LRT for Inferring CPDAG Models
BF-BIC and BF-LRT were developed to support causal search algorithms for estimating nonlinear CPDAG models—that is, Markov equivalence classes of DAGs without latent confounders. In practice, they can serve as plug-in tools wherever a nonlinear score or conditional independence test is required, or both.
@@ 6 Simulation Comparisons
We compare BOSS/BF-BIC, PC-Max/BF-LRT, and PC-Max/KCI (Zhang et al.,, 2012) and PC-Max/RCIT (Strobl et al.,, 2019) in smaller-scale simulations. We use KCI and RCIT from the causal-learn package (Zheng et al.,, 2024). KCI uses kernel reasoning to ensure that $cov(f(X),g(Y))=0$ for all $f$ and $g$ for $X{\;\perp\!\!\!\perp\;}Y$, for the unconditional case, and uses similar reasoning for the conditional case. Note that KCI in causal-learn is fully parallelized, with all matrix calculations optimized for multi-core execution. RCIT chooses random Fourier features to model non-linear relationships. Like BF-BIC, these methods consider functions of $X$ and $Y$, though in the context of the theorem that $cov(f(X),g(Y))$ being zero for all $f$ and $g$ if and only if $X{\;\perp\!\!\!\perp\;}Y$. Fourier features are random sin or cosine functions of $X$ or $Y$. Because these implementations are slower than the BF methods for large $N$ and $p$, we include them only for the continuous 10 node average degree 2 cases. [^3] All simulations are performed on a MacBook Pro M2 with 64 GB of RAM and 10 cores. RCIT scales fairly well though is dominated by the basis function methods. This may be partially due to its being implemented in Python, which is a slower language than Java. For speed, though, an implementation in C would be advantageous, which we have also pursued (https://github.com/bja43/causal-get).
The CAM method proposed by Bühlmann et al., (2014) is directly relevant to our setting, as it assumes a sparse, additive structural equation model with Gaussian noise. [^4] For a feature comparison of CAM, additive noise models, and our BF-BIC and BF-LRT methods, please see Table 1 in the Appendix. Although we had initially planned to include CAM in our benchmarking suite, we encountered technical challenges: the official R package for CAM has been removed from CRAN and is no longer actively maintained. To provide a meaningful point of comparison, however, we replicated the experimental setup from Figure 3 of the CAM paper, generating random DAGs with 10 nodes, average in-degree 2, and 1000 samples, using additive nonlinear functions as in Peters et al., (2014) with Gaussian noise. On this benchmark, BOSS with BF-BIC achieved an average SHD of 5.40 (±3.85), adjacency precision 1.00 (±0.00), and adjacency recall 0.90 (±0.06) over 10 runs. These results indicate strong performance in the same setting used to validate CAM.
@@ 8 A Real Data Example
8.1 Galaxy Analysis
Beyond this environmental application, FCIT and the associated basis-function methods have recently been used in astrophysical research at much larger scales. Desmond and Ramsey, (2025) applied FCIT Ramsey and Andrews, (2025) with both the Basis-Function LRT and Basis-Function BIC to a sample of roughly $5\times 10^{5}$ low-redshift galaxies from the NASA Sloan Atlas, recovering a hierarchical, mass-driven causal structure among galaxy properties while distinguishing these relations from observational selection effects. This study, submitted to MNRAS, demonstrates that our nonlinear score and test framework scales effectively to astronomical sample sizes and yields interpretable causal hypotheses in complex physical systems.
@@ 9 Conclusion
In our real-data analysis of the Algerian Forest Fire dataset, we applied the BOSS (Andrews et al.,, 2023) algorithm and recovered plausible causal structures in the form of Partial Ancestral Graphs (PAGs). In the full model, which included deterministic Canadian Fire System indices, FCIT correctly identified the expected edge $\text{FWI}\rightarrow\text{Fire}$ along with much of the computational structure underlying fire risk indices. The measured-only model also recovered interpretable predictors such as RH and Temperature, reinforcing the reliability of our approach. These results suggest that combining FCIT with BF-based scores and tests yields informative models even in the presence of nonlinearities and latent confounding.
@@
Thanks to Clark Glymour for helpful comments on earlier drafts.
BA was supported by the US National Institutes of Health under the Comorbidity: Substance Use Disorders and Other Psychiatric Conditions Training Program T32DA037183. JR was supported by the US Department of Defense under Contract Number FA8702-15-D-0002 with Carnegie Mellon University for the operation of the Software Engineering Institute. PS was supposed by NIH Award Number: 1043252, “Interpretable graphical models for large multi-model COPD data” and NSF award number 2229881, “AI Institute for Societal Decision Making.” The content of this paper is solely the responsibility of the authors and does not necessarily represent the official views of these funding agencies. The authors have no conflict of interest to report.